OpenStudy (anonymous):

QUESTION 4 These three jobs were in the paper today: Frank's Furters $5.67 /hr. Harriet's Hair Styles $80 for 15 hours weekly. Babysitting $250 a month for 10 hours after school a week. (Assume 4 weeks is a month.) Which job has the highest hourly pay? Frank's Furters Harriet's Hair Styles Babysitting Two of them are tied.

OpenStudy (anonymous):

i think is babysitting

OpenStudy (mynameisnemo):

What's 80 divided by 15

OpenStudy (mynameisnemo):

The answer to that will be the hourly pay for Harriet's Hair Styles.

OpenStudy (anonymous):

Margaret said that 1/5 of her class was absent today. What percent of the class was absent? 10% 15% 20% 30%

OpenStudy (anonymous):

5,3

OpenStudy (mynameisnemo):

Next do $250 divided by four. Once you figure out the answer to that divide that by 10. That will be the hourly pay for Babysitting.

OpenStudy (anonymous):

62.5

OpenStudy (mynameisnemo):

1/5 times 10 on top and bottom = 10/50 multiply that by 2 and you get 20/100. 20% of her class was absent

OpenStudy (mynameisnemo):

and 62.5 seems like a little much for babysitting huh? try again, right numbers but the decimal is in the wrong place.

OpenStudy (anonymous):

how about this one

OpenStudy (anonymous):

Margaret said that 1/5 of her class was absent today. What percent of the class was absent? 10% 15% 20% 30%

OpenStudy (anonymous):

A?

OpenStudy (texaschic101):

1/5 = 0.2 = 20%

OpenStudy (anonymous):

QUESTION 10 Write as a percent: 0.564 5.64% 56.4% 0.564% 564%

OpenStudy (texaschic101):

move the decimal 2 spaces to the right and what do you get ?

OpenStudy (anonymous):

c??

OpenStudy (mynameisnemo):

@lacacakakas I already helped explain the class one

OpenStudy (mynameisnemo):

and @lacacakakas try movin it two spaces the other direction

OpenStudy (texaschic101):

to turn a decimal to percent....either move the decimal 2 spaces to the right....or multiply by 100

OpenStudy (anonymous):

D right

OpenStudy (texaschic101):

not D 0.564....move decimal 2 spaces to the right = 56.4 % or multiply by 100....0.564 * 100 = 56.4 %

OpenStudy (anonymous):

13% of $67.99 =